# Thermodynamic Entropic Uncertainty Relation

### v3

@@ Abstract
Introduction.— The thermodynamic uncertainty relation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] expresses a universal trade-off between the entropy production and the relative variance of time-integrated observables for systems operating far from equilibrium. For a steady-state stochastic thermodynamic system and a trajectory observable $F$, the thermodynamic uncertainty relation states [1, 2]
@@
where $\mathbb{E}[F]$ and $\mathrm{Var}[F]$ are the expectation value and variance of $F$, respectively, and $\Sigma$ denotes the entropy production during the observation interval. The thermodynamic uncertainty relation implies that achieving higher precision inevitably requires greater entropy production–a manifestation of the “no free lunch” principle in thermodynamics. The thermodynamic uncertainty relation given by Eq. (1) quantifies the uncertainty of the observables via their variance; however, uncertainty is often evaluated using the Shannon entropy in the context of information theory. Consider a thermodynamic decision-making process that yields a categorical output, such as a “yes” or “no” label. Since the output is categorical rather than numerical, we cannot use variance to quantify its uncertainty. However, even in such cases, we can still calculate the Shannon entropy. Therefore, to analyze the tradeoff between the accuracy of the categorical outputs and the entropy production, a thermodynamic uncertainty relation based on Shannon entropy is required.
@@
The issue of how to quantify uncertainty has been actively studied in quantum mechanics. In the Heisenberg-Robertson uncertainty relation [11, 12], the uncertainty in the outcomes of two observables is quantified through their variances. Later, Hirschman showed that the variance can be replaced by the Shannon entropy [13, 14]. Maassen and Uffink [15] derived a precise statement of this entropic uncertainty relation [16]: the sum of the entropies of the two observables cannot fall below a bound determined by the extent of their eigenstate overlap. The entropic uncertainty relation plays a fundamental role in quantum cryptography, especially in quantum key distribution protocols [17].
@@
In this Letter, in the context of stochastic thermodynamics, we show that there is an uncertainty relation between entropy production and the symmetry entropy [cf. Eq. (8)], which quantifies the extent of symmetry of the observable’s distribution. Specifically, we show that the sum of the entropy production and the symmetry entropy of the observable should be no less than $\ln 2$ [cf. Eqs. (10) and (14)]. In other words, the asymmetry of the observable probability distribution, as quantified by entropy, requires that the entropy production be at least equal to this measure of asymmetry. As a corollary of the result, we also show that the sum of the entropy production and the Shannon entropy of the observable should be no less than $\ln 2$ [Eq. (12)]. We apply the derived relation to the diffusion decision model, illustrating a trade-off between decision accuracy and entropy production in stochastic decision-making processes. The trade-off relation in thermodynamic cost has been actively studied, particularly in terms of thermodynamic uncertainty relations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] using the variance of observable quantities and speed limits using the distance between states [18, 19, 20, 9, 21, 22, 23]. This study demonstrates a trade-off between the entropy of the observable and the entropy production, which is expected to lead to the derivation of other trade-off relations.
@@
Figure 1: Conceptual representation of the thermodynamic entropic uncertainty relation. (a) Stochastic thermodynamic process. The thermodynamic entropic uncertainty relation considers a stochastic process, where each state transition is random. (b) Trajectory of the stochastic process shown in (a). $\Gamma$ denotes time evolution of a realization of the process. $\Gamma^{\dagger}$ is the time-reversal of $\Gamma$. (c) Probability distribution of observable $\Phi(\Gamma)$. $\Phi(\Gamma)$ is arbitrary, provided it satisfies the time-reversal antisymmetric property [Eq. (7)].
@@
Figure 2: Examples of the symmetry entropy $\Lambda[P(\Phi)]$. Horizontal axes denote values of $\Phi$ (left column) and $|\Phi|$ (right column). The values of $\Phi$ always form pairs; $-1$ and $1$, and $-2$ and $2$ are two pairs in the examples. Vertical axes denote probability distribution $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). (a) Distribution symmetric with respect to $\Phi=0$ where probabilities of all of the pairs are identical. $P(|\Phi|)$ is different from the original distribution $P(\Phi)$ and thus $\Lambda[P(\Phi)]$ is $\ln 2$. (b) Asymmetric distribution where probabilities of all of the pairs are fully biased. $P(|\Phi|)$ and $P(\Phi)$ are identical and thus $\Lambda[P(\Phi)]$ is 00. (c) Asymmetric distribution where probabilities of all of the pairs are fully biased. The distributions $P(|\Phi|)$ and $P(\Phi)$ are different, but they effectively become the same if the labels for the pair $-1$ and $1$ are swapped. Therefore, they are essentially the same distribution. This results in $\Lambda[P(\Phi)]=0$.
Methods.—Let $X$ be a random variable and $P(X)$ be its probability distribution. The Shannon entropy $H[P(X)]$ is defined by
@@
$D[P(X)\|P(Y)]\geq D[P(\tilde{X})\|P(\tilde{Y})],$ (5)
where $P(\tilde{X})$ and $P(\tilde{Y})$ are probability distributions of the transformed variables $\tilde{X}$ and $\tilde{Y}$, respectively.
Having introduced basic concepts of the divergence, we move to consider stochastic thermodynamic systems. Stochastic thermodynamics considers processes whose state change is described by a stochastic process (Fig. 1(a)). Let $\Gamma$ be a stochastic trajectory of the process and $\Gamma^{\dagger}$ be its time reversal (Fig. 1(b)). Moreover, we can define the probability of measuring $\Gamma$, which is denoted by $\mathcal{P}(\Gamma)$. Assuming the local detailed balance, it is known that the entropy production under the steady-state condition is defined by the Kullback-Leibler divergence [24, 25]:
@@
Note that the expression of Eq. (6) also holds for continuous processes such as Langevin dynamics. Consider an observable $\Phi(\Gamma)$, which is a function of the trajectory $\Gamma$. Here, we assume that $\Phi(\Gamma)$ is antisymmetric under the time reversal:
@@
Figure 3: (a) Description of the diffusion decision model. The green trajectory illustrates a realization that exceeds the threshold $\theta$, resulting in a decision of $+1$ (“yes” class). Conversely, the purple trajectory represents a realization that falls below $-\theta$, leading to a decision of $-1$ (“no” class). (b) and (c) Entropies as a function of parameter $f$. In both figures, the blue dashed line shows the entropy production $\Sigma$, and the purple dot-dashed line represents $H[P(\Phi)]$, the entropy of the distribution $P(\Phi)$. Their sum, $\Sigma+H[P(\Phi)]$, is also depicted by the red solid line. The orange dotted line represents $\ln 2$ and $(1-P(0))\ln 2$ in (a) and (b), respectively, which serves as the lower bound for $\Sigma+H[P(\Phi)]$ according to Eq. (15). These plots are shown for $\theta=0$ in (b) and $\theta=0.5$ in (c), with other parameters being $A=1$ and $\tau=10$.
Results.—We derive the thermodynamic entropic uncertainty relation, which is the main result of this Letter. We first introduce an entropic measure which quantifies the observable of a trajectory $\Gamma$. Let $\Lambda[P(\Phi)]$ be
@@
which is the entropy difference between the original distribution $P(\Phi)$ and its absolute valued distribution $P(|\Phi|)$. Here, we call $\Lambda[P(\Phi)]$ symmetry entropy. The symmetry entropy $\Lambda[P(\Phi)]$ quantifies the extent of symmetry of $P(\Phi)$. Intuitively speaking, the difference $H[P(\Phi)]-H[P(|\Phi|)]$ quantifies how much “sign” information remains in the distribution $P(\Phi)$ once its magnitude $P(|\Phi|)$ is known. If $P(\Phi)$ is perfectly symmetric with respect to $\Phi=0$, knowing $P(|\Phi|)$ tells us nothing about the sign, so $\Lambda[P(\Phi)]$ reaches its maximum of $\ln 2$ (one bit). In contrast, if $P(\Phi)$ is fully biased in one direction, the sign is already determined by $P(|\Phi|)$ and $\Lambda[P(\Phi)]$ vanishes. Figure 2 illustrates examples of the symmetry entropy, with the horizontal axes representing the values of $\Phi$ (left column) and $|\Phi|$ (right column), and the vertical axes corresponding to $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). From the time-reversal antisymmetric condition [Eq. (7)], if $\Phi=a$ exists ($a>0$), then $\Phi=-a$ also exists, which is regarded as a pair of the observable. Suppose that the distribution $P(\Phi)$ is symmetric with respect to $\Phi=0$ for all of the pairs, as depicted in Fig. 2(a). For this case, $P(|\Phi|)$ is very different from the original distribution $P(\Phi)$, which results in $\Lambda[P(\Phi)]=\ln 2$. In contrast, for an asymmetric distribution illustrated in Fig. 2(b), $P(\Phi)$ and $P(|\Phi|)$ are the same, and thus $\Lambda[P(\Phi)]$ reduces to 00. Similarly, $\Lambda[P(\Phi)]$ is 00 for the bias pattern represented in Fig. 2(c). When $\Phi(\Gamma)$ is a binary function (outputs two categories), we have $\Lambda[P(\Phi)]=H[P(\Phi)]$.
For simplicity, let us consider the case where the observable $\Phi(\Gamma)$ does not take the value 00, a condition which holds in several problem settings. For instance, we may consider binary classification based on trajectories of stochastic processes. In this case, the observable $\Phi(\Gamma)$ does not include 00. $\Phi(\Gamma)=0$ should be handled separately because $\Phi(\Gamma)=\Phi(\Gamma^{\dagger})=0$ from Eq. (7), showing that the observable is invariant under the time reversal. Later, we will consider the case where $\Phi(\Gamma)=0$ is included. It can be shown that
@@
whose proof is provided in the End Matter. In Eq. (9), $\Lambda[P(\Phi)]$ being 00 and $\ln 2$ corresponds to asymmetric and symmetric distributions, respectively. Using the symmetry entropy $\Lambda[P(\Phi)]$, we obtain the trade-off between the entropy production and the asymmetry of $P(\Phi)$ quantified by entropy:
@@
Equation (10) is the main result of this study and referred to as thermodynamic entropic uncertainty relation. The derivation is shown in the End Matter. The right-hand side of Eq. (10) quantifies the asymmetry of the probability distribution $P(\Phi)$. Equation (10) demonstrates that for any observable $\Phi$ that satisfies the time-reversal antisymmetric condition specified in Eq. (7), the entropy production must be at least $\ln 2-\Lambda[P(\Phi)]$. The trade-off between entropy production and observable asymmetry parallels traditional thermodynamic uncertainty relations, which illustrate a trade-off between the relative variance of observables and entropy production. The thermodynamic uncertainty relations consider the relative variance $\mathrm{Var}[\Phi]/\mathbb{E}[\Phi]^{2}$. In a sense, using the relative variance can also be seen as quantifying the asymmetry of $P(\Phi)$; when the variance is smaller and the expectation is greater, the probability distribution $P(\Phi)$ is more asymmetric with respect to $\Phi=0$. Since $\ln 2-\Lambda[P(\Phi)]\geq 0$ in Eq. (10) due to Eq. (9), Eq. (10) can be regarded as a refinement of the second law, employing the entropy of the observable. There are some advantages in employing the entropy instead of the variance. When the observable is a categorical variable that outputs labels such as “yes” or “no”, which is the case in classification tasks, the use of variance is not appropriate. Even in such cases, the Shannon entropy is well defined, indicating that Eq. (10) can be applied. The right-hand side of Eq. (10) can also be interpreted as a divergence. If we denote the Jensen-Shannon divergence by $\mathrm{JS}[P(X)\|P(Y)]$, then
@@
the derivation of which is shown in the End Matter. Equation (10) provides a trade-off between the symmetry entropy and the entropy production. For a discrete probability distribution, $H[P(|\Phi|)]\geq 0$ holds. Therefore, $H[P(\Phi)]\geq\Lambda[P(\Phi)]$ and thus the following bound also holds:
@@
So far, we assumed that $\Phi(\Gamma)$ takes values in a countable set, that is, $P(\Phi)$ is a discrete distribution. Consider the case where $\Phi(\Gamma)$ produces continuous values, where the summation should be replaced by an integration. When considering a continuous distribution, the notable difference is that the differential entropy may take negative values. However, the existence of negative values is not problematic, as such negative values are offset in $H[P(\Phi)]-H[P(|\Phi|)]$. Therefore, the main result of Eq. (10) is well defined for the continuous case as well.
The derivation of Eq. (10) assumed that the observable $\Phi(\Gamma)$ does not include $\Phi(\Gamma)=0$. It is straightforward to extend the result to the case where the observable includes $\Phi(\Gamma)=0$. Specifically, when $\Phi(\Gamma)$ can equal 00, the range of $\Lambda[P(\Phi)]$ is modified as follows:
$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2$ (13)
@@
which includes Eq. (10) as the specific case $P(\Phi=0)=0$. As long as the entropy production is given by Eq. (6), Eq. (14) holds for any observable $\Phi(\Gamma)$ satisfying Eq. (7). In Eq. (14), $\Phi=0$ plays a special role. When considering the process of doing nothing, the observable $\Phi$ is always 00 implying $P(\Phi=0)=1$. For such empty dynamics, the entropy production is 00, showing that both sides of Eq. (14) equal 00 and thus the inequality becomes an equality. However, note that this exceptional handling of $\Phi=0$ is not limited to Eq. (14). In the conventional thermodynamic uncertainty relation [Eq. (1)], when the expectation of the current vanishes, the inequality is ill-defined. Equations (10) and (14) saturate if and only if the system in equilibrium. For a detailed discussion on the saturation condition, see the supplementary material [26]. Again, when $\Phi(\Gamma)$ is a discrete distribution, the following relation holds as well:
@@
where $\Delta S_{\mathrm{m}}$ is the entropy increase in the surrounding medium. Note that the Landauer principle can be considered as a specific case of the second law of thermodynamics [28]. Given the assumption that the system is in steady state, the change in system entropy $\Delta S$ is zero; consequently, $\Sigma=\Delta S_{\mathrm{m}}$. Then, Eq. (10) can be expressed by $\Delta S_{\mathrm{m}}\geq\ln 2-\Lambda[P(\Phi)]$. When the probability distribution of $P(\Phi)$ ends up with a totally asymmetric distribution, i.e., $\Lambda[P(\Phi)]=0$, Eq. (10) is formally identical to the Landauer principle. However, note that the right side of Eq. (16) arises from a reduction in the Shannon entropy within the system’s state, implying that the Landauer principle is relevant when the system state changes over the time evolution. This contrasts with Eq. (10), where the system is assumed to be in steady state, and $\ln 2$ on its right-hand side arises from the asymmetry of the observable, not the state of the system.
Example.—We demonstrate an application of the thermodynamic entropic uncertainty relation. Decision-making, a fundamental aspect of information processing, is crucial in diverse fields ranging from biology to information science. To illustrate this, we consider the diffusion decision model [29, 30, 31], which describes decision-making as a stochastic process of evidence accumulation. Let $\Gamma$ be a stochastic trajectory of the decision model within the time interval $[0,\tau]$ and $\Phi(\Gamma)$ be a function of $\Gamma$. In the diffusion decision model, a decision occurs when the accumulated evidence during $[0,\tau]$ surpasses thresholds. We assume that if the evidence is greater than or equal to $\theta$, then $\Phi=+1$ (‘yes”‘ class); if it is less than or equal to $-\theta$, then $\Phi=-1$ (“no” class); and in all other cases, $\Phi=0$ (neutral class). Figure 3(a) describes an example of the decision process. Here, $+1,-1,0$ are simply labels, so we can assign other labels. Specifically, let us consider the following Langevin dynamics for modeling the evidence accumulation often employed in the literature [29, 30, 31]:

### v5

@@ Abstract
I Introduction
The thermodynamic uncertainty relation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] expresses a universal trade-off between the entropy production and the relative variance of time-integrated observables for systems operating far from equilibrium. For a steady-state stochastic thermodynamic system and a trajectory observable $F$, the thermodynamic uncertainty relation states [1, 2]
@@ I Introduction
where $\mathbb{E}[F]$ and $\mathrm{Var}[F]$ are the expectation value and variance of $F$, respectively, and $\Sigma$ denotes the entropy production during the observation interval. The thermodynamic uncertainty relation implies that achieving higher precision inevitably requires greater entropy production, a manifestation of the “no free lunch” principle in thermodynamics. The thermodynamic uncertainty relation given by Eq. (1) quantifies the uncertainty of the observables via their variance; however, uncertainty is often evaluated using the Shannon entropy in the context of information theory. Consider a thermodynamic decision-making process that yields a categorical output, such as a “yes” or “no” label. Since the output is categorical rather than numerical, we cannot use variance to quantify its uncertainty. However, even in such cases, we can still calculate the Shannon entropy. Therefore, to analyze the trade-off between the accuracy of the categorical outputs and the entropy production, a thermodynamic uncertainty relation based on Shannon entropy is required.
@@
The issue of how to quantify uncertainty has been actively studied in quantum mechanics. In the Heisenberg-Robertson uncertainty relation [11, 12], the uncertainty in the outcomes of two observables is quantified through their variances. Later, Hirschman showed that the variance can be replaced by the Shannon entropy [13, 14]. Maassen and Uffink [15] derived a precise statement of this entropic uncertainty relation [16]; the sum of the entropies of the two observables cannot fall below a bound determined by the extent of their eigenstate overlap. The entropic uncertainty relation plays a fundamental role in quantum cryptography, especially in quantum key distribution protocols [17].
@@
In this Letter, in the context of stochastic thermodynamics, we show that there is an uncertainty relation between entropy production and the symmetry entropy [cf. Eq. (8)], which quantifies the extent of symmetry of the observable’s distribution. Specifically, we show that the sum of the entropy production and the symmetry entropy of the observable should be no less than $\ln 2$ [cf. Eqs. (10) and (14)]. In other words, the asymmetry of the observable probability distribution, as quantified by entropy, requires that the entropy production be at least equal to this measure of asymmetry. As a corollary of the result, we also show that the sum of the entropy production and the Shannon entropy of the observable should be no less than $\ln 2$ [Eq. (12)]. We apply the derived relation to the diffusion decision model, illustrating a trade-off between decision accuracy and entropy production in stochastic decision-making processes. The trade-off relation in thermodynamic cost has been actively studied, particularly in terms of thermodynamic uncertainty relations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] using the variance of observable quantities and speed limits using the distance between states [18, 19, 20, 9, 21, 22, 23]. This study demonstrates the trade-off between the entropy of the observable and the entropy production, which is expected to lead to the derivation of other trade-off relations.
@@
Figure 1: Conceptual representation of the thermodynamic entropic uncertainty relation. (a) Stochastic thermodynamic process. The thermodynamic entropic uncertainty relation considers a stochastic process, where each state transition is random. (b) Trajectory of the stochastic process shown in (a). $\Gamma$ denotes the time evolution of a realization of the process. $\Gamma^{\dagger}$ is the time-reversal of $\Gamma$. (c) Probability distribution of observable $\Phi(\Gamma)$. $\Phi(\Gamma)$ is arbitrary, provided it satisfies the time-reversal antisymmetric property [Eq. (7)].
@@
Figure 2: Examples of the symmetry entropy $\Lambda[P(\Phi)]$. Horizontal axes denote values of $\Phi$ (left column) and $|\Phi|$ (right column). The values of $\Phi$ always form pairs; 1-1 and $1$, and 2-2 and $2$ are two pairs in the examples. Vertical axes denote the probability distribution $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). (a) Distribution symmetric with respect to $\Phi=0$, where the probabilities of all of the pairs are identical. $P(|\Phi|)$ is different from the original distribution $P(\Phi)$ and thus $\Lambda[P(\Phi)]$ is $\ln 2$. (b) Asymmetric distribution where the probabilities of all of the pairs are fully biased. $P(|\Phi|)$ and $P(\Phi)$ are identical and thus $\Lambda[P(\Phi)]$ is 0. (c) Asymmetric distribution where probabilities of all of the pairs are fully biased. The distributions $P(|\Phi|)$ and $P(\Phi)$ are different, but they effectively become the same if the labels for the pair 1-1 and $1$ are swapped. Therefore, they are essentially the same distribution. This results in $\Lambda[P(\Phi)]=0$.
II Methods
Let $X$ be a random variable and $P(X)$ be its probability distribution. The Shannon entropy $H[P(X)]$ is defined by
@@ II Methods
[…]
where P(X~)P($\tilde${X}) and P(Y~)P($\tilde${Y}) are probability distributions of the transformed variables $\tilde{X}$ and $\tilde{Y}$, respectively.
Having introduced basic concepts of the divergence, we move to consider stochastic thermodynamic systems. Stochastic thermodynamics considers processes whose state changes are described by a stochastic process (Fig. 1(a)). Let $\Gamma$ be a stochastic trajectory of the process and $\Gamma^{\dagger}$ be its time reversal (Fig. 1(b)). Moreover, we can define the probability of measuring $\Gamma$, which is denoted by $\mathcal{P}(\Gamma)$. Assuming the local detailed balance, it is known that the entropy production under the steady-state condition is defined by the Kullback-Leibler divergence [24, 25]:
@@
Note that the expression in Eq. (6) also holds for continuous processes such as Langevin dynamics. Consider an observable $\Phi(\Gamma)$, which is a function of the trajectory $\Gamma$. Here, we assume that $\Phi(\Gamma)$ is antisymmetric under the time reversal:
@@
III Results
We first introduce an entropic measure that quantifies a trajectory observable. Let $\Lambda[P(\Phi)]$ be
@@ III Results
which is the entropy difference between the original distribution $P(\Phi)$ and its absolute-value distribution $P(|\Phi|)$. Here, we call $\Lambda[P(\Phi)]$ symmetry entropy. The symmetry entropy $\Lambda[P(\Phi)]$ quantifies the extent of symmetry of $P(\Phi)$. Intuitively speaking, the difference $H[P(\Phi)]-H[P(|\Phi|)]$ quantifies how much “sign” information remains in the distribution $P(\Phi)$ once its magnitude $P(|\Phi|)$ is known.
For simplicity, let us consider the case where the observable $\Phi(\Gamma)$ does not take the value 0, a condition which holds in several problem settings. For instance, we may consider binary classification based on trajectories of stochastic processes. In this case, the observable $\Phi(\Gamma)$ does not include 0. $\Phi(\Gamma)=0$ should be handled separately because $\Phi(\Gamma)=\Phi(\Gamma^{\dagger})=0$ from Eq. (7), showing that the observable is invariant under the time reversal. Later, we will consider the case where $\Phi(\Gamma)=0$ is included. It can be shown that
@@
whose proof is provided in Appendix B. Let us elaborate on the necessary and sufficient conditions for achieving 0 and $\ln 2$ in Eq. (9). Figure 2 illustrates examples of the symmetry entropy, with the horizontal axes representing the values of $\Phi$ (left column) and $|\Phi|$ (right column), and the vertical axes corresponding to $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). From the time-reversal antisymmetric condition [Eq. (7)], if $\Phi=\phi$ exists ($\phi>0$), then $\Phi=-\phi$ also exists, which is regarded as a pair of the observable. From the calculation in Appendix B, $\Lambda[P(\Phi)]=\ln 2$ if and only if $P(\Phi=\phi)=P(\Phi=-\phi)$ for all $\phi$. This corresponds to the symmetric probability distribution as shown in Fig. 2(a). In contrast, $\Lambda[P(\Phi)]=0$ is satisfied if and only if $P(\Phi=\phi)=0$ or $P(\Phi=-\phi)=0$ for all $\phi$. This case corresponds to asymmetric probability distributions as shown in Figs. 2(b) and (c). When $\Phi(\Gamma)$ is a binary function (outputs two categories), we have $\Lambda[P(\Phi)]=H[P(\Phi)]$.
Using the symmetry entropy $\Lambda[P(\Phi)]$, we obtain the trade-off between the entropy production and the asymmetry of $P(\Phi)$ quantified by entropy:
@@
Equation (10) is the main result of this study and is referred to as the thermodynamic entropic uncertainty relation. The derivation is shown in Appendix A. The right-hand side of Eq. (10) quantifies the asymmetry of the probability distribution $P(\Phi)$. Equation (10) demonstrates that for any observable $\Phi$ that satisfies the time-reversal antisymmetric condition specified in Eq. (7), the entropy production must be at least $\ln 2-\Lambda[P(\Phi)]$. The trade-off between entropy production and observable asymmetry parallels traditional thermodynamic uncertainty relations, which illustrate a trade-off between the relative variance of observables and entropy production. The thermodynamic uncertainty relations consider the relative variance $\mathrm{Var}[\Phi]/\mathbb{E}[\Phi]^{2}$. In a sense, using the relative variance can also be seen as quantifying the asymmetry of $P(\Phi)$; when the variance is smaller and the expectation is greater, the probability distribution $P(\Phi)$ is more asymmetric with respect to $\Phi=0$. Since $\ln 2-\Lambda[P(\Phi)]\geq 0$ in Eq. (10) due to Eq. (9), Eq. (10) can be regarded as a refinement of the second law, employing the entropy of the observable. There are some advantages in employing the entropy instead of the variance. When the observable is a categorical variable that outputs labels such as “yes” or “no”, which is the case in classification tasks, the use of variance is not appropriate. Even in such cases, the symmetry entropy defined by Eq. (8) can handle categorical outputs. As mentioned, the outputs of the observables constitute pairs, i.e., $\Phi=\phi$ and $\Phi=-\phi$. As the value of $\phi$ does not affect the calculation of the entropy $H[P(\Phi)]$, the paired outputs can be replaced with categorical outputs such as $\Phi=\phi=\text{``yes''}$ or $\Phi=-\phi=\text{``no''}$. The right-hand side of Eq. (10) can also be interpreted as a divergence. If we denote the Jensen-Shannon divergence by $\mathrm{JS}[P(X)\|P(Y)]$, then
@@
the derivation of which is shown in Appendix C.
Equation (10) provides a trade-off between the symmetry entropy and the entropy production. For a discrete probability distribution, $H[P(|\Phi|)]\geq 0$ holds. Therefore, $H[P(\Phi)]\geq\Lambda[P(\Phi)]$ and thus the following bound also holds:
@@
So far, we assumed that $\Phi(\Gamma)$ takes values in a countable set, that is, $P(\Phi)$ is a discrete distribution. Consider the case where $\Phi(\Gamma)$ produces continuous values, where the summation should be replaced by an integration. When considering a continuous distribution, the notable difference is that the differential entropy may take negative values. However, the existence of negative values is not problematic, as such negative values are offset in $H[P(\Phi)]-H[P(|\Phi|)]$. Therefore, the main result of Eq. (10) holds for the continuous case as well.
So far, we assumed that the observable $\Phi(\Gamma)$ does not include $\Phi(\Gamma)=0$. It is straightforward to extend the result to the case where the observable includes $\Phi(\Gamma)=0$. Specifically, when $\Phi(\Gamma)$ can equal 0, the range of $\Lambda[P(\Phi)]$ is modified as follows (see Appendix B):
$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2.$ (13)
@@
which includes Eq. (10) as the specific case $P(\Phi=0)=0$. The derivation is shown in Appendix A. As long as the entropy production is given by Eq. (6), Eq. (14) holds for any observable $\Phi(\Gamma)$ satisfying Eq. (7). In Eq. (14), $\Phi=0$ plays a special role. When considering the process of doing nothing, the observable $\Phi$ is always 0, implying $P(\Phi=0)=1$. For such empty dynamics, the entropy production is 0, showing that both sides of Eq. (14) equal 0 and thus the inequality becomes an equality. Although the conventional thermodynamic uncertainty relation [Eq. (1)] does not handle $\Phi=0$ as an exceptional case, it becomes ill-defined for the equilibrium condition where both the expectation of the current and the entropy production vanish. Equations (10) and (14) saturate if and only if the system is in equilibrium. For a detailed discussion on the saturation condition, see Appendix D. Again, when $\Phi(\Gamma)$ is a discrete distribution, the following relation holds as well:
@@
where $\Delta S_{\mathrm{m}}$ is the entropy increase in the surrounding medium. Note that the Landauer principle can be considered as a specific case of the second law of thermodynamics [29]. Given the assumption that the system is in steady state, the change in system entropy $\Delta S$ is zero; consequently, $\Sigma=\Delta S_{\mathrm{m}}$. Then, Eq. (10) can be expressed by $\Delta S_{\mathrm{m}}\geq\ln 2-\Lambda[P(\Phi)]$. When the probability distribution $P(\Phi)$ is totally asymmetric, i.e., $\Lambda[P(\Phi)]=0$, Eq. (10) is formally identical to the Landauer principle. However, note that the right-hand side of Eq. (16) arises from a reduction in the Shannon entropy within the system’s state, implying that the Landauer principle is relevant when the system state changes over the time evolution. This contrasts with Eq. (10), where the system is assumed to be in steady state, and $\ln 2$ on its right-hand side arises from the asymmetry of the observable, not the state of the system.
IV Example
Figure 3: Description of the diffusion decision model. The green trajectory illustrates a realization that exceeds the threshold $\theta$, resulting in a decision of +1+1 (“yes” class). Conversely, the purple trajectory represents a realization that falls below $\theta$, leading to a decision of 1-1 (“no” class). Trajectories that do not fall into either category correspond to a decision of 0 (“neutral” class).
Figure 4: Results for the diffusion decision model with the discrete-output observable [Eq. (18)]. Panels (a) and (b) show entropic quantities as a function of parameter $f$ for $\theta=0$ $\theta=0.5$, respectively, with $A=1$ and $\tau=10$. Panel (c) shows entropic quantities as a function of $\tau$ with $\theta=0$, $A=1$, and $f=0.5$. In these figures, the blue dashed line shows the entropy production $\Sigma$, and the purple dot-dashed line represents $H[P(\Phi)]$. Their sum, $\Sigma+H[P(\Phi)]$, is also depicted by the red solid line. The orange dotted line represents $\ln 2$ in (a) and (c) and $(1-P(0))\ln 2$ in (b), which serves as the lower bound for $\Sigma+H[P(\Phi)]$ according to Eq. (15).
We demonstrate an application of the thermodynamic entropic uncertainty relation. Decision-making, a fundamental aspect of information processing, is crucial in diverse fields ranging from biology to information science. To illustrate this, we consider the diffusion decision model [30, 31, 32], which describes decision-making as a stochastic process of evidence accumulation. Let $\Gamma$ be a stochastic trajectory of the decision model within the time interval $0,\tau]$ and $\Phi(\Gamma)$ be a function of $\Gamma$. In the diffusion decision model, a decision occurs when the accumulated evidence during $0,\tau]$ surpasses thresholds. We assume that if the evidence is greater than or equal to $\theta$, then $\Phi=+1$ (“yes” class); if it is less than or equal to $\theta$, then $\Phi=-1$ (“no” class); and in all other cases, $\Phi=0$ (neutral class). Figure 3 describes an example of the decision process. Here, +1,1,0+1,-1,0 are simply labels, so we can assign other labels. Specifically, let us consider the following Langevin dynamics for modeling the evidence accumulation often employed in the literature [30, 31, 32]: